# Mejdl Safran and Imran Ashraf

### v2

@@ 3.2.3 AutoComplete
Autocomplete feature of the EmoWrite system predicts the word on a character-to-character basis. For this purpose, four different datasets are created, each containing words related to an emotional class. The words are predicted from the respective dataset as per the user’s emotional state.
@@ 3.2.6 Interface Arrangement
(a) First Set of Characters.
(b) Character set after selection of ’t’.
Figure 2: Keyboard Interface.
@@ 3.2.8 Conversion of Thought to Text
To convert thought to text, the mental commands can be detected through the brain signals $I_{s}$ and compared with the given trained signals $T_{s}=\{left,\ right,\ up,\ down\}$. $I{{}_{s_{i}}}(t)$ is the input signal of the $i^{th}$ channel at time $t$ which will be compared with the trained signals.
$\small M\left(t\right)=\sum_{i=1}^{14}\frac{I_{s_{i}}(t)}{T_{s}}$ (1)
@@
$\small B(D_{c})=\left\{\begin{matrix}1&D_{c}\in T_{s}\\ 0&D_{c}\notin T_{s}\\ \end{matrix}\right.$ (3)
@@
$\small S_{L}\left(D_{c}\right)=\left\{\begin{matrix}1&f=\mathrm{Blink},f\in F_{E}\land F_{o}=twice\\ 0&f\neq Blink\\ \end{matrix}\right.$ (4)
@@
$\small{Dis}_{k}=\left|\begin{matrix}\begin{matrix}e&t\\ i&n\\ \end{matrix}&\begin{matrix}a&o\\ \leftarrow&more\\ \end{matrix}\\ \end{matrix}\right|$ (5)
@@
$\small PNextChar\left(x\right)=\prod_{i=1}^{26}{P(x_{i})/P(x)}\ \ \ \ x\equiv S_{L},{\ \ \ \ x}_{i}\in U$ (6)
@@
$\small{Dis}_{e}=\left|\begin{matrix}\begin{matrix}r&d\\ a&t\\ \end{matrix}&\begin{matrix}s&n\\ \leftarrow&more\\ \end{matrix}\\ \end{matrix}\right|$
@@
$\small H_{v}(T)=\left\{\begin{matrix}x&T\in S,\ x\in Sin\\ y&T\in P,\ y\in Plu\\ \end{matrix}\right.$ (7)
@@
$\centering\small\begin{matrix}E_{A}\left(t_{f}-t_{i}\right)\rightarrow high,low\\ E_{V}\left(t_{f}-t_{i}\right)\rightarrow positive,negative\\ D_{E}=E_{A}+E_{V}\end{matrix}\@add@centering$ (8)
@@
$\small P\left(m_{1},.\ .\ .,m_{n}\right)=\prod_{j=1}^{n}{P(m_{j}|m_{1},.\ .\ .,m_{j-1})}$ (9)
@@
$\small s_{t}=\tanh(X_{w_{t}}+Z_{s_{t-1}})$ (10)
$\small O_{t}=softmax(Y_{s_{t}})$ (11)
@@
$\small L\left(t,O\right)=-\frac{1}{W}\sum_{n\in W}{t_{n}logO_{n}}$ (12)
@@
$\small\footnotesize{EmoPred(D_{E},O_{t})}=\left\{\begin{matrix}\begin{matrix}p&D_{E}=happy,p\in(H\cap{\rm RNN}_{W})\\ q&D_{E}=sad,p\in(S\cap{\rm RNN}_{W})\\ \end{matrix}\\ \begin{matrix}r&D_{E}=angry,p\in(A\cap{\rm RNN}_{W})\\ s&D_{E}=calm,p\in(C\cap{\rm RNN}_{W})\\ \end{matrix}\\ \end{matrix}\right.$ (13)
4 Experimentation and Results
@@ 4.1 Keyboard Efficiency
To evaluate the efficiency of the dynamic keyboard, we use two keyboards in the experiment, i.e., (1) a QWERTY keyboard with scan-through keys, and (2) a Dynamic keyboard (EmoWrite). There are two rounds in this experiment. In the first round, the participants are asked to write 10 words (i.e., number, could, who, down, then, which, these, water, long, and about) using each keyboard, one by one. The starting time, when the user starts thinking of the command, and the ending time, when the participants are done with writing the word, is noted and the total time taken is calculated for each word. While in the second round, the participants are asked to write a whole sentence, in this case “the Brain-computer interface helps in processing brain signals”, using each keyboard, and its total time is also calculated. Figure 3 shows the total time taken required to type these words using both keyboards. It is observed that EmoWrite takes less time to type in all the words while the QWERTY keyboard with scan-through keys requires more time to type in all the words.
Figure 3: Word typing speed with QWERTY and EmoWrite
Figure 4: Average time to type a word over 10 trials
@@ 4.3 Information Transfer Rate
Table 4: Comparison of ITR in bits/minutes
@@ 4.4 Accuracy and Latency
To evaluate the accuracy of the EmoWrite system, we determine the difference between the intended target and the observed target. The intended target is the character the participants intend to select while the observed target is the target that is being selected by the brain signals due to an error. For this experiment, the participants are asked to speak loudly about what they are going to write and then the selected target is observed. Each written character is marked as correct or incorrect. In the end, the total number of correctly written characters is calculated. The resulting accuracy of the proposed EmoWrite system is 90.36%.
Furthermore, to check the latency of EmoWrite, we measure the mean and standard deviation of delay (which is the time taken by the user when he/she starts thinking of a command until he/she performs it). The user is shown a certain command (left, right, up, or down), and the time is noted until the user successfully achieved that target. The mean delay of EmoWrite is 2.685 seconds, which is less than the delay of the system proposed by 30 which has a mean delay of 3 seconds.
4.5 System Ease of Use
Further, the system’s ease of use is determined by measuring the stress level of each participant during experimentation with the help of facial expressions. If a participant is frowning while using the system, it indicates that the participant is having some difficulty in using the system. Whenever the stress is increased to a certain limit, a message shows up, indicating an increase in the stress level. For all experimentation, only a few measurements showed the participant to be frowning. Since these measurements are less than 20% of the total measurements, the system is considered not stressful and easy to use.
@@ 4.6 Effectiveness of Integrating Emotion-based Prediction
The novelty of this system is the emotion-based predictions, and to check its effectiveness, we conduct an experiment consisting of two rounds. In the first round, the participants are asked to write some simple sentences without observing their emotional state. The sentences used in the first round are “I am tired of my job”, “I don’t like this world”, “That was awesome”, “I love this world”, and “I have an infection”. In the second round, the participants are asked to write these sentences again but with the integration of their emotional states. For this, first, the participants are shown some video depicting an emotional state. The difference is observed in the word predictions concerning the emotional states i.e., happiness, anger, sadness, and calm. Figure 5 shows the difference in the time required to type the sentences with and without emotional states. It is observed that using emotion-based prediction is more effective as it gives a prediction of emotion-related words like “horribly”, “ughhh”, “awesome” and “terrible”, and this requires less time to type.
Figure 5: Results showing the effectiveness of emotion-based prediction
@@ 5 Conclusion
Funding
This research is funded by the Research Chair of Online Dialogue and Cultural Communication, King Saud University, Riyadh, Saudi Arabia.
Ethical approval
”Not applicable.”
@@
Acknowledgement
The authors extend their appreciation to the Research Chair of Online Dialogue and Cultural Communication, King Saud University, Riyadh, Saudi Arabia for funding this research.

### v3

@@ 3.2.3 AutoComplete
The autocomplete feature of the EmoWrite system predicts the word on a character-to-character basis. For this purpose, four different datasets are created, each containing words related to an emotional class. The words are predicted from the respective dataset as per the user’s emotional state.
@@ 3.2.8 Conversion of Thought to Text
To convert thought to text, the mental commands are detected through the brain signals $I_{s}$ and compared with the given trained signals $T_{s}=\{left,\ right,\ up,\ down\}$. $I{{}_{s_{i}}}(t)$ is the input signal of the $i^{th}$ channel at time $t$ which will be compared with the trained signals.
$M\left(t\right)=\sum_{i=1}^{14}\frac{I_{s_{i}}(t)}{T_{s}}$ (1)
@@
$B(D_{c})=\left\{\begin{matrix}1&D_{c}\in T_{s}\\ 0&D_{c}\notin T_{s}\\ \end{matrix}\right.$ (3)
@@
$S_{L}\left(D_{c}\right)=\left\{\begin{matrix}1&f=\mathrm{Blink},f\in F_{E}\land F_{o}=twice\\ 0&f\neq Blink\\ \end{matrix}\right.$ (4)
@@
${Dis}_{k}=\left|\begin{matrix}\begin{matrix}e&t\\ i&n\\ \end{matrix}&\begin{matrix}a&o\\ \leftarrow&more\\ \end{matrix}\\ \end{matrix}\right|$ (5)
@@
$PNextChar\left(x\right)=\prod_{i=1}^{26}\frac{P(x_{i})}{P(x)}\ \ \ \ x\equiv S_{L},{\ \ \ \ x}_{i}\in U$ (6)
@@
${Dis}_{e}=\left|\begin{matrix}\begin{matrix}r&d\\ a&t\\ \end{matrix}&\begin{matrix}s&n\\ \leftarrow&more\\ \end{matrix}\\ \end{matrix}\right|$
@@
$H_{v}(T)=\left\{\begin{matrix}x&T\in S,\ x\in Sin\\ y&T\in P,\ y\in Plu\\ \end{matrix}\right.$ (7)
Figure 3: Total time required to type the 10 words using QWERTY and EmoWrite keyboards
@@
$\centering\begin{matrix}E_{A}\left(t_{f}-t_{i}\right)\rightarrow high,low\\ E_{V}\left(t_{f}-t_{i}\right)\rightarrow positive,negative\\ D_{E}=E_{A}+E_{V}\end{matrix}\@add@centering$ (8)
@@
$P\left(m_{1},.\ .\ .,m_{n}\right)=\prod_{j=1}^{n}{P(m_{j}|m_{1},.\ .\ .,m_{j-1})}$ (9)
@@
$s_{t}=\tanh(X_{w_{t}}+Z_{s_{t-1}})$ (10)
$O_{t}=softmax(Y_{s_{t}})$ (11)
@@
$L\left(t,O\right)=-\frac{1}{W}\sum_{n\in W}{t_{n}logO_{n}}$ (12)
@@
$\footnotesize{EmoPred(D_{E},O_{t})}=\left\{\begin{matrix}\begin{matrix}p&D_{E}=happy,p\in(H\cap{\rm RNN}_{W})\\ q&D_{E}=sad,p\in(S\cap{\rm RNN}_{W})\\ \end{matrix}\\ \begin{matrix}r&D_{E}=angry,p\in(A\cap{\rm RNN}_{W})\\ s&D_{E}=calm,p\in(C\cap{\rm RNN}_{W})\\ \end{matrix}\\ \end{matrix}\right.$ (13)
4 Results
@@
Figure 4: Average time required to type a word over 10 trials using EmoWrite keyboard
@@ 4.1 Keyboard Efficiency
To evaluate the efficiency of the dynamic keyboard, we use two keyboards in the experiment, i.e., (1) a QWERTY keyboard with scan-through keys, and (2) a Dynamic keyboard (EmoWrite). There are two rounds in this experiment. In the first round, the participants are asked to write 10 words (i.e., number, could, who, down, then, which, these, water, long, and about) using each keyboard, one by one. The starting time, when the user starts thinking of the command, and the ending time, when the participants are done with writing the word, is noted and the total time taken is calculated for each word. While in the second round, the participants are asked to write a whole sentence, in this case “the Brain-computer interface helps in processing brain signals”, using each keyboard, and its total time is also calculated. Figure 3 shows the total time required to type these words using both keyboards. It is observed that EmoWrite takes less time to type in all the words while the QWERTY keyboard with scan-through keys requires more time to type in all the words.
@@ 4.3 Information Transfer Rate
Table 4: Comparison of ITR in bits/minutes of EmoWrite with existing study
@@ 4.4 Accuracy and Latency
To evaluate the accuracy of the EmoWrite system, we determine the difference between the intended target and the observed target. The intended target is the character the participants aim to select while the observed target is the character selected based on brain signals, potentially due to errors. In the experiment, participants are instructed to speak aloud about what they intend to write, and the system records the selected character. Each written character is then marked as either correct or incorrect. The accuracy is computed by calculating the total number of correctly written characters compared to the total number of characters attempted. The resulting accuracy of the proposed EmoWrite system is 90.36%.
Furthermore, to check the latency of EmoWrite, we measure the mean and standard deviation of delay (which is the time taken by the user when he/she starts thinking of a command until he/she performs it). The user is shown a certain command (left, right, up, or down), and the time is noted until the user successfully achieves that target. The mean delay of EmoWrite is 2.685 seconds, which is less than the delay of the system proposed by 30 which has a mean delay of 3 seconds.
4.5 Participant Stress Monitoring during System Interaction
To assess participant responses to the system, we monitored stress levels through facial expressions. If a participant exhibited signs of stress, such as frowning, it indicated potential difficulty with the system. When stress reached a predefined threshold, a notification was displayed. Analysis revealed that fewer than 20% of measurements indicated signs of stress, suggesting the system is generally well-received and does not induce significant stress. This approach provides valuable insights into participant reactions. However, we acknowledge that additional methods could further enhance our understanding of user experience.
@@ 4.6 Effectiveness of Integrating Emotion-based Prediction
The novelty of this system is the emotion-based predictions, and to check its effectiveness, we conduct an experiment consisting of two rounds. In the first round, the participants are asked to write some simple sentences without observing their emotional state. The sentences used in the first round are “I am tired of my job”, “I don’t like this world”, “That was awesome”, “I love this world”, and “I have an infection”. In the second round, the participants are asked to write these sentences again but with the integration of their emotional states. For this, the participants are first shown some video depicting an emotional state. The difference is observed in the word predictions concerning the emotional states i.e., happiness, anger, sadness, and calm. Figure 5 shows the difference in the average time required to type the sentences with and without emotional states. It is observed that using emotion-based prediction is more effective as it gives a prediction of emotion-related words like “horribly”, “ughhh”, “awesome” and “terrible”, and this requires less time to type.
Figure 5: Average time required to type sentences with and without incorporating emotional states.
5 Discussion
This paper introduces EmoWrite, a system designed to aid paralytic individuals in converting their thoughts into text using a sentiment analysis approach. EmoWrite leverages EEG signals, facial expressions, and emotional states to facilitate effective communication. The system has demonstrated promising results in terms of accuracy and typing speed. However, the study acknowledges certain limitations and areas for improvement:
5.1 Training Time
One of the primary limitations is the extensive training time required for participants to effectively use EmoWrite. Each user must undergo a substantial training period to achieve a high level of proficiency with the system. This training involves learning to control the interface with mental commands and facial expressions, as well as adapting to the personalized predictive text system.
5.2 Failure Cases
The analysis of failure cases in this study is based on the accuracy of the results. Instances where the system makes incorrect predictions are evaluated to understand the underlying causes and improve the system’s robustness. These errors can be attributed to factors such as signal noise, user fatigue, and variations in individual brain signal patterns.
5.3 Generalizability
While EmoWrite has been tailored to suit individual users by incorporating their emotional states and typing patterns, its generalizability to a broader population with varying types of paralysis and cognitive conditions remains an area for further research.
5.4 User Interface
The dynamic and adaptive nature of the virtual keyboard in EmoWrite is a significant advancement, but the interface’s complexity can pose a challenge to new users. Simplifying the interface without compromising functionality could enhance user experience and reduce the learning curve.
5.5 Future Improvements
Future iterations of EmoWrite could benefit from several enhancements to further its effectiveness. Incorporating advanced machine learning algorithms could improve the system’s accuracy and reduce prediction errors. Additionally, exploring alternative EEG headsets with higher signal resolution and reduced noise may enhance the system’s overall performance and reliability.
Moreover, a direct measure of cognitive load should be integrated into the system’s evaluation to provide a more comprehensive understanding of usability. Improvements in user training protocols and refining the system’s accuracy and usability will also be key focus areas.
In conclusion, EmoWrite represents a significant advancement in assistive communication technologies for individuals with severe disabilities. While the current version demonstrates innovative potential through its integration of sentiment analysis with thought-to-text conversion, addressing its limitations through ongoing research and development will be crucial in maximizing its impact and enhancing the quality of life for users.
@@ 6 Conclusion
Ethics Declarations
“Prior to study initiation, written informed consent was obtained from all participants and the study protocol was approved by the Ethics Committee of COMSATS University Islamabad, Lahore Campus. All experiments were performed in accordance with relevant guidelines and regulations. Informed consent was obtained from all participants for publication of identifying information/images in an online open-access publication.”